# Spatial Proportional Hazards Model with Differential Regularization

### v1

@@ 4.1 Characterization of the Solution to the Penalized Partial Likelihood Problem
We assume that the design matrix $XX^{\top}$ is of full rank and that the weight matrix $W$ contains strictly positive entries. Define $H=X(X^{\top}WX)^{-1}X^{\top}W$ and $Q=I-H$, where $I$ is an identity matrix of the appropriate size. For any function $u$ within the functional space $\mathcal{F}=H^{2}_{n}(\Omega)$, let $u_{n}=(u(p_{1}),\ldots,u(p_{n}))$ denote the vector of evaluations of $u$ at the $n$ spatial locations $\{p_{i}\}_{i=1}^{n}$. Let $\tilde{\beta}$ and $\tilde{f}$ denote the minimizers of the functional $\tilde{J}_{\lambda}(\beta,f)$ in (6), and $\hat{\beta}$ and $\hat{f}$ the maximizers of the penalized partial likelihood functional $\mathcal{L}_{p}(\beta,f)$ in (4). Under these conditions, the following Proposition describes the minimizers $\tilde{\beta}$ and $\tilde{f}$ of the functional (6).
@@
There exists a unique pair $(\tilde{\beta},\tilde{f})\in\mathbb{R}^{q}\times H^{2}_{n}$ that minimizes (4). Furthermore,
  • $\tilde{\beta}=(X^{\top}WX)^{-1}X^{\top}W(z-\tilde{f}_{n})$, where $\tilde{f}_{n}=(\tilde{f}(p_{1}),\ldots,\tilde{f}(p_{n}))^{\top}$,
@@
    $u_{n}^{\top}Q\tilde{f}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=u_{n}^{\top}Qz,\quad\forall u\in H^{2}_{n}.$ (7)
@@ 4.2 Solution to the Penalized Partial Likelihood
According to Wilhelm and Sangalli (2016); Sangalli et al. (2013), problem (7) is equivalent to finding $(\tilde{f},\tilde{h})\in H^{1}_{n}(\Omega)\times H^{1}_{n}(\Omega)$ such that
$\begin{cases}u_{n}^{\top}Q\tilde{f}_{n}-\lambda\int_{\Omega}(\nabla u)^{\top}\nabla\tilde{h}=u_{n}^{\top}Qz,\\ -\int_{\Omega}(\nabla\tilde{f})^{\top}\nabla v=\int_{\Omega}\tilde{h}v,\end{cases}$ (8)
@@ 4.3 Finite Elements
The finite element method is a prevalent numerical technique in engineering for solving problems involving partial differential equations (Quarteroni and Quarteroni, 2009).
To build a finite element space, we begin by dividing the domain of interest $\Omega$ into smaller subdomains. Common domain partitions include triangular meshes. We consider a regular triangulation $\mathcal{T}$ of $\Omega$, where adjacent triangles share either a vertex or an edge. The domain $\Omega$ is thus approximated by $\Omega_{\mathcal{T}}$, which is the union of all triangles. Consequently, the boundary $\partial\Omega$ is approximated by a polygon (or multiple polygons for domains with interior holes). This triangulation effectively captures the complex geometry of the domain, including significant concavities such as rivers and detailed structures like bridges that connect different parts of the city.
From this triangulation, locally supported polynomial functions are defined over the triangles, forming a set of basis functions $\psi_{1},\ldots,\psi_{K}$ that span a finite-dimensional subspace $\mathcal{F}_{K}$ of $H^{1}_{n}$. Linear finite elements are created using a basis system where each basis function $\psi_{i}$ is associated with a vertex $\xi_{i}$ (for $i=1,\ldots,K$) of the triangulation $\mathcal{T}$. Each basis function $\psi_{i}$ is a piecewise linear polynomial that equals one at the vertex $\xi_{i}$ and zero at all other vertices of the mesh, i.e., $\psi_{i}(\xi_{j})=\delta_{ij}$ for all $i,j=1,\ldots,K$, where $\delta_{ij}$ is the Kronecker delta. Figure 1 provides an example of a linear finite element basis function on a planar mesh, highlighting its localized support.
Let $\psi=(\psi_{1},\ldots,\psi_{K})^{\top}$ be the column vector containing the $K$ piecewise linear basis functions corresponding to the $K$ vertices $\xi_{1},\ldots,\xi_{K}$. Any function $h$ in the finite element space $\mathcal{F}_{K}$ can be expressed as a linear combination of the basis functions $\psi_{1},\ldots,\psi_{K}$. Let $h=(h_{1},\ldots,h_{K})$ represent the coefficients of this basis expansion for $h$, such that
@@ 4.4 Numerical solution to the Penalized Partial Likelihood
The functions and integrals in equation (8) can be approximated using elements from the finite element space $\mathcal{F}_{K}$. As a result, the problem (8) is approximated by its discrete version: find $(\tilde{f},\tilde{h})\in\mathcal{F}_{K}\times\mathcal{F}_{K}$ such that (8) holds for all $(u,v)\in\mathcal{F}_{K}\times\mathcal{F}_{K}$, where the integrals are now computed over the triangulation $\Omega_{\mathcal{T}}$. Define $\Psi$ as the $n\times K$ matrix representing the evaluations of the $K$ basis functions at the $n$ data locations $\mathbf{p}_{1},\ldots,\mathbf{p}_{n}$,
$\Psi=\begin{bmatrix}\psi^{\top}(\mathbf{p}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{p}_{n})\end{bmatrix}$ (9)
@@
$\mathbf{R}_{0}=\int_{\Omega_{\mathcal{T}}}(\psi\psi^{\top}),\quad\mathbf{R}_{1}=\int_{\Omega_{\mathcal{T}}}(\nabla\psi)^{\top}\nabla\psi.$
@@
$\int_{\Omega_{\mathcal{T}}}(\nabla u)^{\top}\nabla\tilde{h}=\mathbf{u}^{\top}\mathbf{R}_{1}\tilde{\mathbf{h}},\quad\int_{\Omega_{\mathcal{T}}}(\nabla\tilde{f})^{\top}\nabla v=\tilde{\mathbf{f}}^{\top}\mathbf{R}_{1}\mathbf{v},\quad\int_{\Omega_{\mathcal{T}}}\tilde{h}\,v=\tilde{\mathbf{h}}^{\top}\mathbf{R}_{0}\mathbf{v},$
@@
$\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\Psi&\lambda\mathbf{R}_{1}\\ \lambda\mathbf{R}_{1}&\lambda\mathbf{R}_{0}\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{f}}\\ \tilde{\mathbf{h}}\end{bmatrix}=\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\mathbf{z}\\ 0\end{bmatrix},$ (10)
@@
The uniqueness of the solution to (10) is guaranteed by the positive definiteness of the matrices $\mathbf{R}_{0}$ and $(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1})$.
Let $\mathbf{P}=\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1}$ and $\mathbf{S}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}$. Using the functional form of the PIRLS algorithm, and leveraging Propositions 1 and 2, we obtain the following expressions for the maximizers $\hat{\bm{\beta}}$ and $\hat{f}$ of the penalized partial likelihood (2):
$\hat{\bm{\beta}}=(\mathbf{X}^{\top}\mathbf{W}\mathbf{X})^{-1}\mathbf{X}^{\top}\mathbf{W}(\mathbf{I}-\mathbf{S})\mathbf{z},$
@@
$\hat{\mathbf{f}}=(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z},$
and the vector of evaluations of $\hat{f}$ at the $n$ data locations given by
$\hat{\mathbf{f}}_{n}=\Psi\hat{\mathbf{f}}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z}=\mathbf{S}\mathbf{z},$
where the vector of pseudo-data $\mathbf{z}$, and the matrices $\mathbf{W},\mathbf{Q}$, and $\mathbf{S}$ are those obtained upon convergence of the PIRLS algorithm.
The positive definite matrix $\mathbf{P}$ represents the discretization of the penalty term in (2) and (7). Due to the variational formulation (8) of the estimation problem, this penalty matrix does not require the computation of second-order derivatives. Azzimonti et al. (2014) demonstrates that in the finite element space used for discretization, $\mathbf{P}$ is equivalent to the penalty matrix that would result from a direct discretization of the penalty term in (2) and (7), which would otherwise involve the computation of second-order derivatives.
@@ 4.5 Selecting the Parameter $\mu$ Using Cross-Validation
When it comes to choosing the parameter $\mu$, one straightforward approach is a visual inspection. This method allows the researcher to select a value that achieves the desired smoothness of the estimated spatial field. However, to avoid subjective decision-making, we propose a more objective, numerical method. Following the approach of Cygu et al. (2021), we can utilize cross-validation to determine the optimal value of $\mu$. Specifically, we select the $\mu$ that results in the lowest cross-validated partial likelihood deviance (CV-PLD) or the highest cross-validated Harrell’s concordance index (CV-C-index) (Harrell Jr et al., 1996).
To compute the CV-PLD for each $\mu$, we employ $k$-fold cross-validation. This process involves splitting the training data into $k$ subsets (folds), training the model on $k-1$ folds, and validating it on the remaining fold. This procedure is repeated $k$ times, each time with a different fold serving as the validation set. We use two metrics for this purpose: CV-PLD and CV-C-index (Dai and Breheny, 2019). The CV-PLD is defined as:
@@
where $\mathcal{L}(\hat{\beta}_{-k}(\mu))$ represents the log partial likelihood evaluated at $\hat{\beta}_{-k}$ using the entire dataset, and $\mathcal{L}_{-k}(\hat{\beta}_{-k}(\mu))$ is the log partial likelihood evaluated at $\hat{\beta}_{-k}$ on the retained data (all but the excluded fold). The $\hat{\beta}_{-k}$ values denote the penalized estimates derived from the retained data. The optimal value of $\mu$ is the one that minimizes (11). Notably, equation (11) typically provides superior results compared to simply evaluating the partial likelihood on the held-out fold (a method sometimes referred to as the basic approach), as the likelihood of any given observation is influenced by the other elements in the risk set.
The alternative cross-validation metric, the CV-C-index, is based on Harrell’s concordance index for Cox models. This metric, known as the cross-validated C-index, calculates the probability that, for a random pair of individuals, the predicted survival times are correctly ordered according to their actual survival times. This method is similar to the implementation available in the survival package (Therneau et al., 2015).

### v2

@@ 4.1 Characterization of the Solution to the Penalized Partial Likelihood Problem
We assume that the design matrix $XX^{\top}$ is of full rank and that the weight matrix $W$ contains strictly positive entries. Define $H=X(X^{\top}WX)^{-1}X^{\top}W$ and $Q=I-H$, where $I$ is an identity matrix of the appropriate size. For any function $u$ within the functional space $\mathcal{F}=H^{2}_{n}(\Omega)$, let $\bm{u}_{n}=(u(\bm{p}_{1}),\ldots,u(\bm{p}_{n}))^{\top}$ denote the vector of evaluations of $u$ at the $n$ spatial locations $\{\bm{p}_{i}\}_{i=1}^{n}$. Let $\tilde{\bm{\beta}}$ and $\tilde{f}$ denote the minimizers of the functional $\tilde{J}_{\lambda}(\beta,f)$ in (6), and $\hat{\beta}$ and $\hat{f}$ the maximizers of the penalized partial likelihood functional $\mathcal{L}_{p}(\bm{\beta},f)$ in (4). Under these conditions, the following proposition describes the minimizers $\tilde{\bm{\beta}}$ and $\tilde{f}$ of the functional (6).
@@
There exists a unique pair $(\tilde{\bm{\beta}},\tilde{f})\in\mathbb{R}^{q}\times H^{2}_{n}$ that minimizes (4). Furthermore,
  • $\tilde{\bm{\beta}}=(X^{\top}WX)^{-1}X^{\top}W(z-\tilde{f}_{n})$, where $\tilde{\bm{f}}_{n}=(\tilde{f}(\bm{p}_{1}),\ldots,\tilde{f}(\bm{p}_{n}))^{\top}$,
@@
    $\bm{u}_{n}^{\top}Q\bm{\tilde{}}{f}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=\bm{u}_{n}^{\top}Qz,\quad\forall u\in H^{2}_{n}.$ (7)
@@ 4.2 Solution to the Penalized Partial Likelihood
According to Wilhelm and Sangalli, (2016); Sangalli et al., (2013), problem (7) is equivalent to finding $(\tilde{f},\tilde{h})\in H^{1}_{n}(\Omega)\times H^{1}_{n}(\Omega)$ such that
$\begin{cases}\bm{u}_{n}^{\top}Q\bm{\tilde{}}{f}_{n}-\lambda\int_{\Omega}(\nabla u)^{\top}\nabla\tilde{h}=\bm{u}_{n}^{\top}Qz,\\ -\int_{\Omega}(\nabla\tilde{f})^{\top}\nabla v=\int_{\Omega}\tilde{h}v,\end{cases}$ (8)
@@ 4.3 Finite Elements
The finite element method is a prevalent numerical technique in engineering for solving problems involving partial differential equations (Quarteroni and Quarteroni,, 2009). To build a finite element space, we begin by dividing the domain of interest $\Omega$ into smaller subdomains. Common domain partitions include triangular meshes. We consider a regular triangulation $\mathcal{T}$ of $\Omega$, where adjacent triangles share either a vertex or an edge. The domain $\Omega$ is thus approximated by $\Omega_{\mathcal{T}}$, which is the union of all triangles. Consequently, the boundary $\partial\Omega$ is approximated by a polygon (or multiple polygons for domains with interior holes). This triangulation effectively captures the complex geometry of the domain, including significant concavities such as rivers and detailed structures like bridges that connect different parts of the city.
From this triangulation, locally supported polynomial functions are defined over the triangles, forming a set of basis functions $\psi_{1},\ldots,\psi_{K}$ that span a finite-dimensional subspace $\mathcal{F}_{K}$ of $H^{1}_{n}$. Linear finite elements are created using a basis system where each basis function $\psi_{i}$ is associated with a vertex $\xi_{i}$ (for $i=1,\ldots,K$) of the triangulation $\mathcal{T}$. In particular, each basis function $\psi_{i}$ is a piecewise linear polynomial that equals one at the vertex $\xi_{i}$ and zero at all other vertices of the mesh, i.e., $\psi_{i}(\xi_{j})=\delta_{ij}$ for all $i,j=1,\ldots,K$, where $\delta_{ij}$ is the Kronecker delta. Figure 1 provides an example of a linear finite element basis function on a planar mesh, highlighting its localized support.
Let $\bm{\psi}=(\psi_{1},\ldots,\psi_{K})^{\top}$ be the column vector containing the $K$ piecewise linear basis functions corresponding to the $K$ vertices $\xi_{1},\ldots,\xi_{K}$. Any function $h$ in the finite element space $\mathcal{F}_{K}$ can be expressed as a linear combination of the basis functions $\psi_{1},\ldots,\psi_{K}$. Let $\bm{h}=(h_{1},\ldots,h_{K})$ represent the coefficients of this basis expansion for $\bm{h}$, such that
@@ 4.4 Numerical solution to the Penalized Partial Likelihood
The functions and integrals in equation (8) can be approximated using elements from the finite element space $\mathcal{F}_{K}$. As a result, the problem (8) is approximated by its discrete version: find $(\tilde{f},\tilde{h})\in\mathcal{F}_{K}\times\mathcal{F}_{K}$ such that (8) holds for all $(u,v)\in\mathcal{F}_{K}\times\mathcal{F}_{K}$, where the integrals are now computed over the triangulation $\Omega_{\mathcal{T}}$. Define $\Psi$ as the $n\times K$ matrix representing the evaluations of the $K$ basis functions at the $n$ data locations $\bm{p}_{1},\ldots,\bm{p}_{n}$,
$\Psi=\begin{bmatrix}\psi^{\top}(\bm{p}_{1})\\ \vdots\\ \psi^{\top}(\bm{p}_{n})\end{bmatrix}$ (9)
@@
$R_{0}=\int_{\Omega_{\mathcal{T}}}(\psi\psi^{\top}),\quad R_{1}=\int_{\Omega_{\mathcal{T}}}(\nabla\psi)^{\top}\nabla\psi.$
@@
$\int_{\Omega_{\mathcal{T}}}(\nabla u)^{\top}\nabla\tilde{h}=\mathbf{u}^{\top}R_{1}\tilde{\mathbf{h}},\quad\int_{\Omega_{\mathcal{T}}}(\nabla\tilde{f})^{\top}\nabla v=\tilde{\mathbf{f}}^{\top}R_{1}\mathbf{v},\quad\int_{\Omega_{\mathcal{T}}}\tilde{h}\,v=\tilde{\mathbf{h}}^{\top}R_{0}\mathbf{v},$
@@
$\begin{bmatrix}-\Psi^{\top}Q\Psi&\lambda R_{1}\\ \lambda R_{1}&\lambda R_{0}\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{f}}\\ \tilde{\mathbf{h}}\end{bmatrix}=\begin{bmatrix}-\Psi^{\top}Q\mathbf{z}\\ \bm{0}\end{bmatrix},$ (10)
@@
Proof.
The uniqueness of the solution to (10) is guaranteed by the positive definiteness of the matrices $R_{0}$ and $(\Psi^{\top}Q\Psi+\lambda R_{1}R_{0}^{-1}R_{1})$. ∎
Let $P=R_{1}R_{0}^{-1}R_{1}$ and $S=\Psi(\Psi^{\top}Q\Psi+\lambda P)^{-1}\Psi^{\top}Q$. Using the functional form of the PIRLS algorithm, and leveraging Propositions 1 and 2, we obtain the following expressions for the maximizers $\hat{\bm{\beta}}$ and $\hat{f}$ of the penalized partial likelihood (2):
$\hat{\bm{\beta}}=(X^{\top}WX)^{-1}X^{\top}W(I-S)\bm{z},$
@@
$\hat{\bm{f}}=(\Psi^{\top}Q\Psi+\lambda P)^{-1}\Psi^{\top}Q\mathbf{z},$
and the vector of evaluations of $\hat{\bm{f}}$ at the $n$ data locations given by
$\hat{\bm{f}}_{n}=\Psi\hat{\bm{f}}=\Psi(\Psi^{\top}Q\Psi+\lambda P)^{-1}\Psi^{\top}Q\bm{z}=S\bm{z},$
where the vector of pseudo-data $\bm{z}$, and the matrices $W,Q$, and $S$ are those obtained upon convergence of the PIRLS algorithm.
The positive definite matrix $P$ represents the discretization of the penalty term in (2) and (7). Due to the variational formulation (8) of the estimation problem, this penalty matrix does not require the computation of second-order derivatives. Azzimonti et al., (2014) demonstrates that in the finite element space used for discretization, $P$ is equivalent to the penalty matrix that would result from a direct discretization of the penalty term in (2) and (7), which would otherwise involve the computation of second-order derivatives.
@@ 4.5 Selecting the Parameter $\mu$ Using Cross-Validation
When it comes to choosing the parameter $\mu$, one straightforward approach is a visual inspection. This method allows the researcher to select a value that achieves the desired smoothness of the estimated spatial field. However, to avoid subjective decision-making, we propose a more objective, numerical method. Following the approach of Cygu et al., (2021), we can utilize cross-validation to determine the optimal value of $\mu$. Specifically, we select the $\mu$ that results in the lowest cross-validated partial likelihood deviance (CV-PLD) or the highest cross-validated Harrell’s concordance index (CV-C-index) (Harrell Jr et al.,, 1996).
To compute the CV-PLD for each $\mu$, we employ $k$-fold cross-validation. This process involves splitting the training data into $k$ subsets (folds), training the model on $k-1$ folds, and validating it on the remaining fold. This procedure is repeated $k$ times, each time with a different fold serving as the validation set. We use two metrics for this purpose: CV-PLD and CV-C-index (Dai and Breheny,, 2019). The CV-PLD is defined as:
@@
where $\mathcal{L}(\hat{\bm{\beta}}_{-k}(\mu),\hat{f}_{-k}(\mu))$ represents the log partial likelihood evaluated at $(\hat{\bm{\beta}}_{-k},\hat{f}_{-k}(\mu))$ using the entire dataset, and $\mathcal{L}_{-k}(\hat{\bm{\beta}}_{-k}(\mu),\hat{f}_{-k}(\mu))$ is the log partial likelihood evaluated at $(\hat{\bm{\beta}}_{-k},\hat{f}_{-k}(\mu))$ on the retained data (all but the excluded fold). The $(\hat{\bm{\beta}}_{-k},\hat{f}_{-k}(\mu))$ values denote the penalized estimates derived from the retained data. The optimal value of $\mu$ is the one that minimizes (11). Notably, equation (11) typically provides superior results compared to simply evaluating the partial likelihood on the held-out fold (a method sometimes referred to as the basic approach), as the likelihood of any given observation is influenced by the other elements in the risk set.
The alternative cross-validation metric, the CV-C-index, is based on Harrell’s concordance index for Cox models. This metric, known as the cross-validated C-index, calculates the probability that, for a random pair of individuals, the predicted survival times are correctly ordered according to their actual survival times. This method is similar to the implementation available in the survival package (Therneau et al.,, 2015).
@@ 5 Model Version for Areal data
6 Simulation Results
To demonstrate the performance of the proposed model, we present simulations on a horseshoe domain, as defined in Ramsay, (2002); Wood et al., (2008). We use the spatial field $f(x,y)$, depicted in Figure 2(a) and defined as follows:
where $r=0.5$ and $r_{0}=0.1$, and $c$ is a constant chosen so that the minimum value of $f$ over the horseshoe domain is zero, approximately $c\simeq 1.77$. Figure]2(a) shows the values of the spatial field over the horseshoe.
6.1 Geostatistical Data
We randomly sample $n=100,200,300$ points, denoted by $\{\bm{p}_{i}\}_{i=1}^{n}$, from the horseshoe domain, as illustrated in Figure 2(b). Figure 2(c) shows the corresponding Delaunay triangulation (Shewchuk,, 2002). Each sampled point is associated with an observation $x_{i}$ of a unidimensional random variable $X_{i}$, for which we omit the bold notation, and that follows a standard normal distribution. Additionally, we observe $(y_{i},\delta_{i})$, where $y_{i}=\min(T_{i},C_{i})$ and $\delta_{i}=I(T_{i}\leq C_{i})$. The variable $T_{i}$ given $X_{i}=x$ follows a proportional hazards model with baseline hazard $\lambda_{0}(t)=1$ and $\beta_{0}=0.2$, as defined in equation (1), with the form $\lambda(t|x)=\lambda_{0}(t)\exp(x\beta_{0}+f(p_{i}))$. This implies that the random variable $T_{i}$ at location ±$\pm p_{i}$ follows an exponential distribution with rate (inverse of the mean) $\exp(x\beta_{0}+f(p_{i}))$.
To introduce censoring, we set $C_{i}=1$ for each $i$, resulting in approximately 20% censored observations. We set $\mu=0.5,1,5,10$ in order to understand how the value of $\mu$ affects the estimation of the spatial field $f$ and of the regression coefficients vector of $\beta$. The finite elements method is applied using a basis of order $1$ constructed on the Delaunay triangulation over the sampled locations. We conduct $N=250$ simulations for each combination of $n$ and $\mu$. An example of the estimated spatial field is given in Figure 2(d).
(a) Spatial field $f(\bm{p})$ over the horseshoe domain
(b) Example of sample points over the horseshoe domain for $n=300$.
(c) Mesh based on the sampled points of Figure 2(b).
(d) Example of estimation of the spatial field based on the sampled points of Figure 2(b) and setting $\mu=10$.
We report in Table 1 the simulation results that highlights how changes in the smoothing parameter $\mu$, censoring levels, and sample sizes affect key performance metrics. the bias values for $\beta$ remain close to zero across most scenarios, indicating unbiased estimates, with slight variations at higher levels of $\mu$. The standard deviation (Sd) decreases as the sample size $n$ increases, suggesting that larger samples improve estimate stability. Additionally, the 95% confidence interval coverage probability (CP95) stays close to the nominal level, particularly in larger samples, supporting the reliability of interval estimates.
For the spatial field $f$, we reported the Mean Squared Error (MSE)defined as $\text{MSE}=n^{-1}\sum_{i=1}^{n}(\hat{f}(\bm{p}_{i})-f(\bm{p}_{i}))^{2}$, as a performance metric. we observe that the MSE values generally decrease as $n$ grows, suggesting improved precision with more data. Higher censoring percentages (e.g., 40%) introduce a slight increase in MSE and its standard deviation (Sd MSE), indicating reduced accuracy under increased censoring.
Regression coefficient $\beta$ Spatial field $f$
$\mu$ Cens. % $n$ bias Sd CP95 MSE Sd MSE
0.5 20% 100 -0.001 0.121 0.944 0.177 0.125
200 -0.010 0.078 0.964 0.155 0.111
40% 100 -0.002 0.145 0.944 0.227 0.222
200 -0.013 0.091 0.960 0.150 0.096
1 20% 100 -0.001 0.122 0.944 0.172 0.126
200 -0.010 0.077 0.968 0.157 0.121
40% 100 -0.003 0.145 0.932 0.206 0.171
200 -0.012 0.091 0.960 0.153 0.110
3 20% 100 -0.001 0.121 0.948 0.177 0.146
200 -0.009 0.077 0.972 0.158 0.137
40% 100 -0.003 0.145 0.936 0.206 0.178
200 -0.012 0.091 0.964 0.155 0.128
5 20% 100 -0.001 0.121 0.948 0.185 0.164
200 -0.009 0.077 0.972 0.161 0.146
40% 100 -0.003 0.145 0.940 0.216 0.196
200 -0.011 0.090 0.964 0.159 0.141
10 20% 100 -0.002 0.121 0.948 0.196 0.177
200 -0.008 0.077 0.972 0.166 0.163
40% 100 -0.002 0.145 0.940 0.229 0.217
200 -0.010 0.090 0.964 0.163 0.157
Table 1: Simulation results for estimating the regression coefficient $\beta$ and spatial field $f$ across different values of the smoothing parameter $\mu$, varying censoring percentages (Cens. %) and sample sizes $n$. The performance metrics include bias, standard deviation (Sd), and 95% confidence interval coverage probability (CP95) for $\beta$, as well as mean squared error (MSE) and standard deviation of MSE (Sd MSE) for $f$. The results demonstrate the impact of $\mu$, censoring level, and sample size on the precision and accuracy of the estimates.
6.2 Areal Data